On June 30, Year 7, Princess Company issued $4,000,000 face value of 13%, 20-year bonds at $4,300,920, a yield of 12%. Princess uses the effective-interest method to amortize bond premiums and discounts. The bonds pay interest semiannually on June 30 and December 31.

Instructions:

Round all answers to the nearest dollar!

A. Prepare the journal entries to record the following transactions:

  1. The issuance of the bonds on June 30, Year 7
  2. The payment of interest and the amortization of the premium on December 31, Year 7
  3. The payment of interest and the amortization of the premium on June 30, Year 8
  4. The payment of interest and the amortization of the premium on December 31, Year 8

B. Show the proper balance sheet presentation for the liability for bonds payable on the

     December 31, Year 8 balance sheet

C. Provide answers to the following questions:

  1. What amount of interest expense is reported for Year 8?
  2. Will the bond interest expense reported in Year 8 be the same as, greater than, or less than

the amount that would be reported if the straight-line method of amortization were used?

  1. Determine the total cost of borrowing over the life of the bonds.
  2. Will the total bond interest expense for the life of the bonds be greater than, the same as, or less than the total interest expense if the straight-line method of amortization were used?
Answer #1
A.) Date Account Titles Debit $ Credit $
June.30 Year 7 Cash 4,300,920
Bond Payable 4,000,000
Premium on Bond    300,920
December 31,Year 7 Interest Expense ( 4,300,920 x 12% x 1/2 )    258,055
Premium on Bond          1,945
Cash ( 4,000,000 x 13% x 1/2 )    260,000
June.30 Year 8 Interest Expense ( 4,300,920 - 1,945 ) x 12% x 1/2    257,939
Premium on Bond          2,062
Cash ( 4,000,000 x 13% x 1/2 )    260,000
December 31,Year 8 Interest Expense ( 4,300,920 - 1,945 - 2,062 ) x 6%    257,815
Premium on Bond          2,185
Cash ( 4,000,000 x 13% x 1/2 )    260,000
B.) Balance Sheet
December 31,2018 Amount $
Non-Current Liabilities
Bond Payable ( Face Value ) 4,000,000
Add : Premium on Bond     294,728
(300,920 - 1,945 - 2,062 - 2,185 )
Bond Carrying Value 4,294,728
C.) Interest Expense for Year 8     515,754 =257939+257815
Straight Line amortization
Interest Expense     504,954 =260000+260000-(300920/20)

The bond interest expense reported in Year 8 will be greater than than the amount that would be reported if the straight-line method

of amortization were used.

Total Cost of Borrowings over the life of Bond $ 10,099,080 =(4000000*13%*20)-300920
The total bond interest expense for the life of the bonds will be the same as the total interest expense if the straight-line method of amortization were used.